You intend to estimate a population proportion with a confidence interval. The data suggests that the normal distribution is a reasonable approximation for the binomial distribution in this case.

While it is an uncommon confidence level, find the critical value that corresponds to a confidence level of 81.2%.
(Report answer accurate to three decimal places with appropriate rounding.)

za/2 = ±
Answer #1

solution

At 81.2%confidence level the z is ,

= 1 - 0.812 = 0.188

/ 2 = 0.188 / 2 = 0.094

Z/2 = Z0.094= ±1.317